OpenStudy (anonymous):

4cos(theta)=-4sqrt(3) sin(-theta) I am supposed to solve the equation on the interval o(

sam (.sam.):

\[4 \cos (x)=-4 \sqrt{3} \sin (-x)\] Can be written as \[4 \cos(x)=4 \sqrt{3} \sin(x) \]

sam (.sam.):

\[4 \cos(x)=4 \sqrt{3} \sin(x)\] \[\frac{4}{4\sqrt{3}}=\tan(x)\] \[\frac{1}{\sqrt{3}}=\tan(x)\]  Find the value of tangent \[x=\tan^{-1}(\frac{1}{\sqrt{3}})=30~degrees\] Then use the chart, since its positive you look at the positive side, which is 180+x So, the other x value is \[x=180+30=210 ~ degrees\]

sam (.sam.):

Converting it to radians \[\frac{30}{180} \times \pi =\frac{1}{6}\pi\] \[\frac{210}{180} \times \pi = \frac{7}{6} \pi\]